OpenStudy (calculusxy):

Solve the equation for 'x' 9(x-5)=4x-5 A. 8 B. -10 C. -8 D. 10

OpenStudy (anonymous):

use distrubte property with 9(x-5) so it's like mutipliying 9x - 45 = 4x -5 +45 to on right side then subtract 4x from 9x so you x by itself 5x = 40 / 5 then you get x =

OpenStudy (anonymous):

\[9x-45=4x-5\] \[-4x+45-4x+45\] \[5x=40\] \[x=8\] A